OpenStudy (anonymous):

What is the slope the line with the equation 5x-2y=10

OpenStudy (anonymous):

you need to rewrite the equation in slope intercept form y = mx+ b

OpenStudy (anonymous):

once you have match it to that form then it is just simply the coefficient of x. And @timo86m it would actually be solve for y

OpenStudy (anonymous):

oh right. solve for y :) my mind is melting too hot here.

OpenStudy (anonymous):

anyway guyes can you guyes please help me to solve

OpenStudy (anonymous):

PLEASE

OpenStudy (anonymous):

first solve for y: -2y=10-5x (divide by -2) y= 5/2(x)-5 this is y=mx+b where M=slope=5/2

OpenStudy (anonymous):

Thank you